# Spin-dependent Force from an NV center Ensemble on a Microlever

### v1

@@
Abstract
We report the observation of spin-dependent force induced by Nitrogen Vacancy (NV) centers embedded in a diamond crystal attached to a tethered oscillator. This result was obtained using a spin-dependent torque generated by a micro-diamond containing billions of NV centers, placed at the end of a commercially available silicon cantilever. We demonstrate that the fundamental flexural mode of the cantilever is excited by the spins of the NV centers. These spins are uniformly coupled to an externally applied homogeneous magnetic field, which enables the NV centers to exert a measurable spin-dependent torque on the cantilever, driving its center of mass motion. Our experiment highlights the intricate interplay between the spin properties of NV centers and the mechanical behavior of the oscillator in the presence of a pure torque, providing a set of tools for exploring quantum effects in state of the art mechanical systems.
@@ Abstract
Our findings provide a new pathway for exploring quantum-to-classical coupling in hybrid systems. In particular, it shows that the fundamental flexural motion of a singly clamped low-Q cantilever can be excited by NV spins without magnetic field gradients. The latter generally bring about magnetic noise Rugar et al. (2004); Kolkowitz et al. (2012) and to not lend themselves easily to the use of many spins. Combined with the low temperature capability, such a platform offers prospects for studying phase transitions such as proposed in Wei et al. Wei et al. (2015) and to scalably couple many spins through the mechanical degree of freedom. We also present a method for mitigating spurious microwave noise, providing valuable insights for research on the mechanical detection of the spin of single centers.
@@ I.1 Mechanical mode properties and detection set-up
The mechanical resonator that we consider in this study is the first bending mode of a commercially available Si cantilever from MikroMasch (HQ:CSC38/tipless/No Al). A picture of the system is shown in Fig. 1-(a). The dimensions of the cantilever are $350~{}\mu{\rm m}\times 32.5~{}\mu{\rm m}\times 1~{}\mu{\rm m}$. We attach a $40~{}\mu{\rm m}$ diameter diamond bought from Adamas nanotechnology (MDNV40umHi30mg) at the end of the cantilever (see Appendix A for explanations on the attachment process). The cantilever, hereafter referred to as ’loaded,’ is enclosed in a vacuum chamber to minimize acoustic noise. However, all experiments presented here were conducted at ambient pressure.
@@
In order to detect the mechanical resonance of the loaded cantilever, we make use of a fibered interferometer. The optical setup is shown in Fig. 1-(b). We use a $638~{}{\rm nm}$ diode laser that is coupled to an optical fiber and sent to a fiber coupler. One of the outputs is directed to a cleaved and stripped fiber positioned in front of the cantilever. The diamond is attached to the other side of the cantilever. The light that is reflected off the fiber end (about 15 % of the incoming power) interferes with the light that is reflected off the cantilever (about 20 %). The cantilever is fixed to three axes Attocube stages (ANSxyz50 model) in order to position the cantilever apex precisely in front of the fiber and to scan it along the optical axis. The light is detected at one of the fiber coupled port by an APD (APD430A from Thorlabs). The signal is divided into three parts: one part is sent to a PID in order to control the $x$ position of the PZT to remain at half fringe during the measurement. The two other parts of the electric signal are sent to a lock-in amplifier and to a spectrum analyzer.
@@
The resonance frequency of the first bending mode was measured prior to attaching the diamond by recording the power spectral density (PSD) on a spectrum analyzer, with no external excitations other than collisions with the surrounding gas. We found this (unloaded) under-damped frequency mode to be at a frequency $f_{0}=14.48~{}{\rm kHz}$, which is consistent with the technical data. The effective mass of the cantilever for this first bending mode can be estimated using the relation $m_{\rm e}\approx\frac{m}{4}$ Rabe et al. (1996), where $m$ is the total mass of the lever. From the cantilever volume and the density of silicon, we find $m_{\rm e}\approx 10^{-11}~{}{\rm kg}$.
@@
The calibrated power spectral density of the signal at APD 1 is shown in Fig. 1-(c). The resonance frequency of this mode is shifted down to $f_{\rm m}=2.856~{}{\rm kHz}$, much lower than the frequency in the unloaded case. The mass of a 40 $\mu$m diamond is on the order of $m_{\rm diam}~{}\approx 10^{-10}~{}{\rm kg}$, which is ten times larger than the effective mass of the first bending mode. The glue also adds weight to the cantilever. One can estimate the total added weight using the formula $f_{\rm m}=\sqrt{k_{\rm m}/(m_{\rm e}+m_{\rm add})}/2\pi$ for the resonance frequency $f_{\rm m}$ of the first loaded bending mode, where $m_{\rm add}$ is the additional mass of both the glue and the diamond. We can estimate the total mass added at the end of the cantilever using the formula:
@@
We find that $m_{\rm add}\approx 25~{}m_{\rm e}$, which is in the right order of magnitude for of the mass of the diamond and the glue.
Fitting the PSD using the theoretical mechanical response to background gas collisions in the presence of laser noise gives us a quality factor of $Q=160$ and a stiffness of $k_{\rm m}\approx 0.03$ N/m. To ensure that this measured mode still corresponds to the first bending mode, we perform a mode profile analysis by moving the cantilever along its main axis while recording the thermal displacement for different positions of the cantilever (see Appendix B). The measured displacement follows the expected trend of the fundamental flexural mode. The thermal displacement is the largest at the apex of the lever and reaches 0.4 nm. This value closely approximates the displacement predicted by the equipartition theorem, assuming the thermal bath is primarily dominated by collisions with a background gas at 300 K. We can thus estimate the sensitivity of the force detection via the cantilever at this temperature. It reads
$F_{\rm min}=\sqrt{\frac{4k_{\rm B}T}{\pi}\frac{k_{\rm m}}{Qf_{\rm m}}}\approx 10^{-14}~{}{\rm N/\sqrt{\rm Hz}}.$ (2)
@@ I.2 NV centers properties and optical set-up
The unit cell of diamond containing a nitrogen-vacancy (NV) center is illustrated in the inset of Fig. 1-(a). The ground electronic state of the negatively charged NV center is a triplet state, characterized by a zero-field splitting that establishes a natural quantization axis, denoted as $\bm{e}_{3}$, along the N-V direction Doherty et al. (2013). In the absence of an external magnetic field, the $|S=1,m_{s}=\pm 1\rangle$ states in the triplet manifold have a frequency $D\approx(2\pi)2.87$ GHz above the $|m_{s}=0\rangle$ state in the ground state at room temperature. Crucially, the electronic spin of the NV center can be optically polarized to the |$\ket{S=1,m_{s}=0}$ state using green laser light. Additionally, the Stokes-shifted photoluminescence (PL) is stronger in the |$\ket{m_{s}=0}$ state compared to the |$\ket{m_{s}=\pm 1}$ states. Consequently, applying a microwave signal at the resonant frequency induces a drop in PL, which forms the basis of optically detected magnetic resonance (ODMR). In the following, the eigenstates |$\ket{m_{s}=i}$ will simply be denoted as |$\ket{i}$.
The NV detection employs the confocal spectroscopy setup shown in Fig. 1-(b), enabling green laser excitation and red photoluminescence collection using a single lens. In the following, the laser power is limited to $1~{}{\rm mW}$ when focused on the entire diamond. At higher laser powers, we observed diamond heating and slow drifts in the mechanical frequency. To detect spin transitions in the electronic spin ground state, we use continuous-wave ODMR, with microwave excitation provided by a Rohde and Schwarz SMB100A signal generator. As will be discussed next, to detect the spin action on mechanical modes, the frequency of the microwave signal will be frequency modulated at the mechanical frequency $f_{m}$.
The diamond at the end of the cantilever is highly doped in NV center with an NV concentration about $3.5~{}{\rm ppm}$. Notably, there are four possible orientations of the N-V axis in the diamond lattice, as the nitrogen atom can occupy any carbon position adjacent to the vacancy. These four orientations, or classes, result in four distinct projections of the magnetic field along the NV axis. When a magnetic field is applied, the degeneracy between these four classes is lifted, typically revealing eight distinct spin transitions instead of two. Given the NV concentration, we found the number of NV centers for a single class to be $N\approx 5\times 10^{9}$.
@@ II Origin of the coupling
In this section, we explain the origin of the mechanical coupling between NV centers and the first bending mode of the cantilever in a presence of a homogeneous magnetic field.
@@ II.1 Cantilever-based torque detection
In his book Foundation of Nanomechanics Cleland (2013), A.N. Cleland examines the scenario in which a cantilever is subjected to a constant torque applied perpendicular to its long axis at its tip. It is shown that this not only causes torsion at the end of the cantilever but also induces bending when the finite thickness of the cantilever is taken into account. This treatment shows that bending modes are sensitive to pure torques and can, therefore, be used for torque detection. In recent years, measuring the frequency shift of the first bending mode due to magnetic torques from anisotropic magnetic materials attached to the end of a nanobeam has become a common technique for studying the magnetic properties of these materials Mehlin et al. (2015); Gross et al. (2016, 2020, 2021).
By restricting our study in the Fourier domain to dynamical motions that are close to the frequency of the first bending mode $f_{\rm m}$, we can neglect the contribution of other mechanical modes in the displacement of the end of the cantilever $x(t)$. It allows us to write the relation $x(t)=l_{\rm e}\phi(t)$ between the angular deflection $\phi(t)$ at the end of the cantilever and $x(t)$, where $l_{\rm e}\approx 190~{}\mu{\rm m}$ is the effective length of the first bending mode. The latter is given by the length of the cantilever divided by the mode shape factor $\beta=1.875$. A torque $\tau_{y}$ along $y$ is applied at the end of the cantilever by the NV centers can thus be converted into a force along $\mathbf{e}_{x}$ via the relation $F_{x}(t)=\tau_{y}(t)/l_{\rm e}$ Gross et al. (2016).
@@
$\tau_{\rm min}\approx 10^{-18}~{}{\rm N.m/\sqrt{\rm Hz}},$ (3)
@@ II.2 Torques induced by NV centers
The spin-spin interaction between the two electrons in the ground state manifold introduces an anisotropy energy term that lifts the degeneracy between the magnetic eigenstates |±$\ket{\pm 1}$ and the non-magnetic eigenstate |$\ket{0}$. This anisotropy term defines a preferential axis for the magnetic moment in the eigenstates $\bm{\mu}_{i}=\mu_{\rm B}\bra{i}\hat{\mathbf{S}}\ket{i}$ to align with. In the presence of an external homogeneous magnetic field $\mathbf{B}_{0}$, the orientation of $\bm{\mu}_{i}$ results from a competition between both the crystalline anisotropy, and the magnetic field’s orientation and strength (see Appendix C for details). If the magnetic field and the crystalline anisotropy (the N-V axis) are not aligned, then $\bm{\mu}_{i}$ will not be aligned with $\mathbf{B}_{0}$, resulting in a net magnetic torque $\bm{\mu}_{i}\times\mathbf{B}_{0}$ applied to the diamond. Finally, the total magnetic torque exerted on the diamond crystal is given by:
@@
where $\bm{\mu}=\sum\rho_{ii}\bm{\mu}_{i}$, with $\rho_{ii}$ being the population in the |$\ket{i}$ state. This torque has been employed to realize spin-cooling of the libration of levitated diamonds embedded with NV centers Delord et al. (2020). Due to the conservation of angular momentum, any changes in the magnetic moment (proportional to angular momentum) of the NV centers also lead to a change in the macroscopic angular momentum of the diamond crystal. This phenomenon is known as the Einstein-de Haas effect Einstein and De Haas (1915), and it is responsible for a torque that has notably been observed with yttrium iron garnet (YIG) disks placed on a torsional resonator Mori et al. (2020). Closely related to this effect, the absorption of microwave photons carrying angular momentum at a magnetic resonance also leads to a net torque on the mechanics. This phenomenon was first observed using a torsional mode of a fiber, where DPPH containing spin-half paramagnetic impurities was placed at the end Alzetta et al. (1967); Arimondo (1968). For further details on calculations of the NV center torques, see Appendix D. As we will show, only the magnetic torque of the form $\bm{\mu}\times\mathbf{B}_{0}$ is in line with our experimental observations.
Figure 3: (a) Optically Detected Magnetic Resonance obtained by measuring the photoluminescence from an ensemble of NV centers. The NV orientations are indicated at the top. Bottom curve: Mechanical detection of magnetic resonance using a frequency modulated microwave (FM-MDMR). (b) (i) and (ii): zoom on the FM-MDMR obtained for the same NV orientations for the |$\ket{0^{\prime}}\to\ket{-1^{\prime}}$ transitions in (i) and for the |$\ket{0^{\prime}}\to\ket{+1^{\prime}}$ transitions in (ii).

### v2

@@
Spin-dependent Force from an NV center Ensemble on a Microlever
Maxime Perdriat Laboratoire De Physique de l’École Normale Supérieure, ENS, PSL, CNRS, Sorbonne Université, Université de Paris, 24 rue Lhomond, 75005 Paris, France.    Alrik Durand1    Louis Chambard1    Julien Voisin1    Gabriel Hétet1
Résumé
We report the observation of spin-dependent force induced by Nitrogen Vacancy (NV) centers embedded in a diamond crystal attached to a tethered oscillator. This result was obtained using a spin-dependent torque generated by a micro-diamond containing billions of NV centers, placed at the end of a commercially available silicon cantilever. Our experiment highlights the intricate interplay between the spin properties of NV centers and the mechanical behavior of the oscillator in the presence of a pure torque, providing a set of tools for exploring quantum effects in state of the art micro-mechanical systems.
@@
Our findings provide a new pathway for exploring quantum-to-classical coupling in hybrid systems. In particular, it shows that the fundamental flexural motion of a singly clamped low-Q cantilever can be excited by NV spins without magnetic field gradients. The latter generally bring about magnetic noise Rugar et al. (2004); Kolkowitz et al. (2012) and to not lend themselves easily to the use of many spins. We also present a method for mitigating spurious microwave noise, providing valuable insights for research on the mechanical detection of the spin of single centers. Combined with the low temperature capability, such a platform offers prospects for studying phase transitions such as proposed in Wei et al. Wei et al. (2015) and to scalably couple many spins through the mechanical degree of freedom.
@@ I.1 Mechanical mode properties and detection set-up
The mechanical resonator that we consider in this study is the first bending mode of a commercially available Si cantilever from MikroMasch (HQ:CSC38/tipless/No Al). A picture of the system is shown in Fig. 1-(a). The dimensions of the cantilever are $350\leavevmode\nobreak\ \mu{\rm m}\times 32.5\leavevmode\nobreak\ \mu{\rm m}\times 1\leavevmode\nobreak\ \mu{\rm m}$. We attach a $40\leavevmode\nobreak\ \mu{\rm m}$ diameter diamond bought from Adamas nanotechnology (MDNV40umHi30mg) at the end of the cantilever (see Appendix A for explanations on the attachment process). The cantilever, hereafter referred to as ’loaded,’ is enclosed in a vacuum chamber to minimize acoustic noise. However, all experiments presented here were conducted at ambient pressure.
@@
In order to detect the mechanical resonance of the loaded cantilever, we make use of a fiber interferometer. The optical setup is shown in Fig. 1-(b). We use a $638\leavevmode\nobreak\ {\rm nm}$ diode laser that is coupled to an optical fiber and sent to a fiber coupler. One of the outputs of the coupler is directed to a cleaved and stripped fiber positioned in front of the cantilever. The diamond is attached to the other side of the cantilever. The cantilever is fixed to three axes Attocube stages (ANSxyz50 model) in order to position its apex precisely in front of the fiber and to scan it along the optical axis. The light that is reflected off the fiber end (about 15 % of the incoming power) then interferes with the light that is reflected off the cantilever (about 20 %). The retro-reflected light is detected at the other fiber coupler port by an APD (APD430A from Thorlabs). This electronic signal is then divided into three parts: one part is sent to a PID in order to control the $x$ position of the PZT to remain at half fringe during the measurement. The two other parts of the electric signal are sent to a lock-in amplifier and to a spectrum analyzer.
@@
The resonance frequency of the first bending mode was measured prior to attaching the diamond by recording the power spectral density (PSD) on a spectrum analyzer, with no external excitations other than collisions with the surrounding gas. We found this (unloaded) under-damped frequency mode to be at a frequency $f_{0}=14.48\leavevmode\nobreak\ {\rm kHz}$, which is consistent with the technical data. The effective mass of the cantilever for this first bending mode can be estimated using the relation $m_{\rm e}\approx\frac{m}{4}$ Rabe et al. (1996), where $m$ is the total mass of the lever. From the cantilever volume and the density of silicon, we find $m_{\rm e}\approx 10^{-11}\leavevmode\nobreak\ {\rm kg}$.
@@
The calibrated power spectral density of the signal at APD 1 is shown in Fig. 1-(c). The resonance frequency of this mode is shifted down to $f_{\rm m}=2.86\leavevmode\nobreak\ {\rm kHz}$. The mass of a 40 $\mu$m diamond is of the order of $m_{\rm diam}\leavevmode\nobreak\ \approx 10^{-10}\leavevmode\nobreak\ {\rm kg}$, which is ten times larger than the effective mass of the first bending mode. The glue also adds mass to the loaded cantilever. One can estimate the total added mass using the formula $f_{\rm m}=\sqrt{k_{\rm m}/(m_{\rm e}+m_{\rm add})}/2\pi$ for the resonance frequency $f_{\rm m}$ of the first loaded bending mode, where $m_{\rm add}$ is the additional mass of both the glue and the diamond. We can estimate $m_{\rm add}$ using the formula:
@@
We find that $m_{\rm add}\approx 25\leavevmode\nobreak\ m_{\rm e}$, which is of the right order of magnitude for the mass of the diamond and the glue.
Fitting the PSD using the theoretical mechanical response to background gas collisions in the presence of laser noise gives us a quality factor of $Q=160$ and a stiffness of $k_{\rm m}\approx 0.03$ N/m. To ensure that this measured mode still corresponds to the first bending mode, we perform a mode profile analysis by moving the cantilever along $x$ while recording the thermal displacement for different positions of the cantilever (see Appendix B). The measured displacement follows the expected trend of the fundamental flexural mode. The thermal displacement is the largest at the apex of the lever and reaches 0.4 nm. This value closely approximates the displacement predicted by the equipartition theorem, assuming the thermal bath is primarily dominated by collisions with a background gas at 300 K. We can thus estimate the sensitivity of the force detection via the cantilever at this temperature. It reads
$F_{\rm min}=\sqrt{\frac{4k_{\rm B}T}{\pi}\frac{k_{\rm m}}{Qf_{\rm m}}}\approx 10^{-14}\leavevmode\nobreak\ {\rm N/\sqrt{\rm Hz}}.$ (2)
@@ I.2 NV centers properties and optical set-up
The unit cell of diamond containing a nitrogen-vacancy (NV) center is illustrated in the inset of Fig. 1-(a). The ground electronic state of the negatively charged NV center is a triplet state, characterized by a zero-field splitting that establishes a natural quantization axis, denoted as $\bm{e}_{3}$, along the N-V direction Doherty et al. (2013). In the absence of an external magnetic field, the $|S=1,m_{s}=\pm 1\rangle$ states in the triplet manifold have a frequency $D\approx(2\pi)2.87$ GHz above the $|m_{s}=0\rangle$ state in the ground state at room temperature. Crucially, the electronic spin of the NV center can be optically polarized to the $|{S=1,m_{s}=0}\rangle$ state using green laser light. Additionally, the Stokes-shifted photoluminescence (PL) is stronger in the $|{m_{s}=0}\rangle$ state compared to the $|{m_{s}=\pm 1}\rangle$ states. Consequently, applying a microwave signal at the resonant frequency induces a drop in PL, which forms the basis of optically detected magnetic resonance (ODMR). In the following, the eigenstates $|{m_{s}=i}\rangle$ will simply be denoted as $|{i}\rangle$.
The diamond at the end of the cantilever is highly doped with NV centers, with an NV concentration of about $3.5\leavevmode\nobreak\ {\rm ppm}$. Notably, there are four possible orientations of the NV axis in the diamond lattice, as the nitrogen atom can occupy any carbon site adjacent to the vacancy. These four orientations, or classes, yield four distinct projections of the magnetic field along the NV axis. When a magnetic field is applied, the degeneracy between these four classes is lifted, typically revealing eight distinct spin transitions instead of two. Given the NV concentration, we estimate the number of NV centers in a single class to be $N\approx 5\times 10^{9}$.
@@ II Origin of the coupling
In this section, we explain the origin of the mechanical coupling between NV centers and the first bending mode of the cantilever in the presence of a homogeneous magnetic field.
@@ II.1 Cantilever-based torque detection
In his book “Foundations of Nanomechanics" Cleland (2013), A. N. Cleland examines the scenario in which a cantilever is subjected to a constant torque applied perpendicular to its long axis at its tip. It is shown that this not only causes torsion at the end of the cantilever but also induces bending when the finite thickness of the cantilever is taken into account. This treatment demonstrates that bending modes are sensitive to pure torques and can, therefore, be used for torque detection. In recent years, measuring the frequency shift of the first bending mode due to magnetic torques from anisotropic magnetic materials attached to the end of a nanobeam has become a common technique for studying the magnetic properties of these materials Mehlin et al. (2015); Gross et al. (2016, 2020, 2021).
By restricting our study in the Fourier domain to dynamical motions that are close to the frequency of the first bending mode $f_{\rm m}$, we can neglect the contribution of other mechanical modes in the displacement of the end of the cantilever $x(t)$. It allows us to write the relation $x(t)=l_{\rm e}\phi(t)$ between the angular deflection $\phi(t)$ at the end of the cantilever and $x(t)$, where $l_{\rm e}\approx 190\leavevmode\nobreak\ \mu{\rm m}$ is the effective length of the first bending mode. The latter is given by the length of the cantilever divided by the mode shape factor $\beta=1.875$. A torque $\tau_{y}$ along $y$ is applied at the end of the cantilever by the NV centers can thus be converted into a force along $\mathbf{e}_{x}$ via the relation $F_{x}(t)=\tau_{y}(t)/l_{\rm e}$ Gross et al. (2016).
@@
$\tau_{\rm min}\approx 10^{-18}\leavevmode\nobreak\ {\rm N\cdot m/\sqrt{\rm Hz}},$ (3)
@@ II.2 Torques induced by NV centers
The spin-spin interaction between the two electrons in the ground state manifold introduces an anisotropy energy term that lifts the degeneracy between the magnetic eigenstates |±$|{\pm 1}\rangle$ and the non-magnetic eigenstate $|{0}\rangle$. This anisotropy term defines a preferential axis for the magnetic moment in the eigenstates $\bm{\mu}_{i}=\mu_{\rm B}\langle{i}|\hat{\mathbf{S}}|{i}\rangle$ to align with. In the presence of an external homogeneous magnetic field $\mathbf{B}_{0}$, the orientation of $\bm{\mu}_{i}$ results from a competition between both the crystalline anisotropy, and the magnetic field’s orientation and strength (see Appendix C for details). If the magnetic field and the crystalline anisotropy (the N-V axis) are not aligned, then $\bm{\mu}_{i}$ will not be aligned with $\mathbf{B}_{0}$, resulting in a net magnetic torque $\bm{\mu}_{i}\times\mathbf{B}_{0}$ applied to the diamond. Finally, the total magnetic torque exerted on the diamond crystal is given by:
@@
where $\bm{\mu}=\sum\rho_{ii}\bm{\mu}_{i}$, with $\rho_{ii}$ being the population in the $|{i}\rangle$ state. This torque has been employed to realize spin-cooling of the libration of levitated diamonds embedded with NV centers Delord et al. (2020). Due to the conservation of angular momentum, any changes in the magnetic moment (proportional to angular momentum) of the NV centers also lead to a change in the macroscopic angular momentum of the diamond crystal. This phenomenon is known as the Einstein-de Haas effect Einstein and De Haas (1915), and it is responsible for a torque that has notably been observed with yttrium iron garnet (YIG) disks placed on a torsional resonator Mori et al. (2020). Closely related to this effect, the absorption of microwave photons carrying angular momentum at a magnetic resonance also leads to a net torque on the mechanics. This phenomenon was first observed using a torsional mode of a fiber, where a DPPH crystal (2,2-diphenyl-1-picrylhydrazyl) containing spin one-half paramagnetic impurities was placed at the end Alzetta et al. (1967); Arimondo (1968). For further details on calculations of the NV center torques, see Appendix D. As we will show, only the magnetic torque of the form $\bm{\mu}\times\mathbf{B}_{0}$ is in line with our experimental observations.
Figure 3: (a) Optically Detected Magnetic Resonance obtained by measuring the photoluminescence from an ensemble of NV centers. The NV orientations are indicated at the top. Bottom curve: Mechanical detection of magnetic resonance using a frequency modulated microwave (FM-MDMR). (b) (i) and (ii): zoom on the FM-MDMR obtained for the same NV orientations for the $|{0^{\prime}}\rangle\to|{-1^{\prime}}\rangle$ transitions in (i) and for the $|{0^{\prime}}\rangle\to|{+1^{\prime}}\rangle$ transitions in (ii).